OpenStudy (anonymous):

The length of a rectangle is 12 in. more than its width. The perimeter is 64 in. What is the length of the rectangle? in.

OpenStudy (anonymous):

Answer 2 times the length plus 2 times the width equals perimeter therefore, 2L+2W= P the length is three times the width therefore, L = 3W and you know the perimeter is 96 and inserting 3w for length you get, 2(3W)+2W=96, then 6W+2W=96, then 8W=96. Divide both sides by 8 and you get W=12. The width is 12. And the length is 36.

OpenStudy (mathstudent55):

@babygirl55555 Which problem are you answering?

OpenStudy (anonymous):

i need help

OpenStudy (mathstudent55):

@EvaMarire Let's call the length L and the width W. The problem states: "The length of a rectangle is 12 in. more than its width." Since the width is W, what is L equal to in terms of the width?

OpenStudy (mathstudent55):

The width is W. What do you do to W to get the length?

OpenStudy (anonymous):

divide

OpenStudy (mathstudent55):

No. Use the sentence form the problem. The sentence is" "the length is 12 in. more than the width." If the width is W, what is 12 more than W?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

is it 24

OpenStudy (mathstudent55):

Exactly. For example: 5 more than 2 is 2 + 5 10 more than 6 is 6 + 10 8 more than x is x + 8 12 more than W is ____________

OpenStudy (anonymous):

but 12 + 24= 72

OpenStudy (mathstudent55):

We don't know what W is. We must keep the width as W. What is 12 more than W? W + 12

OpenStudy (mathstudent55):

Not quite. 12 + 24 = 36, not 72

OpenStudy (mathstudent55):

Ok, do you understand what I wrote above? The length is 12 more than the width. Since the width is W, the length is W + 12.

OpenStudy (mathstudent55):

Now we know this: width: W length: W + 12

OpenStudy (anonymous):

so the answer 24

OpenStudy (mathstudent55):

Right. That tells you what the perimeter is. The perimeter, is P. P = 2(L + W) We have L = W + 12, so we get P = 2(W + 12 + W) We are also told the perimeter is 64, so we can replace P with 64: 64 = 2(2W + 12) Now we divide both sides by 2: 32 = 2W + 12 Now we subtract 12 from both sides: 20 = 2W Divide both sides by 2: 10 = W The width is 10. The length is 12 more than the width, so the length is 10 + 12 = 22

OpenStudy (anonymous):

so its 24?

OpenStudy (mathstudent55):

The length is 22 in.

OpenStudy (mathstudent55):

We can check. The width is 10 in. 12 in. more than 10 in. is 22 in. P = 2(22 in. + 10 in.) = 2(32 in.) = 64 in. The problem does state that the perimeter is 64 in., so our answer of length = 22 in. is correct.

OpenStudy (anonymous):

its not 20 l=20

OpenStudy (mathstudent55):

No. Length = 22 in.

OpenStudy (anonymous):

can u help with one more

OpenStudy (anonymous):

What is the value of x?

OpenStudy (anonymous):

can u help me

OpenStudy (mathstudent55):

Ok.

OpenStudy (mathstudent55):

Here is the equation: \(\dfrac{x}{5} + 6 = 36\)

OpenStudy (mathstudent55):

You want to solve for x. That means you want x by itself. You need to see what operations are being done to x and you need to undo those operations.

OpenStudy (mathstudent55):

The fraction x/5 means x is being divided by 5. Then "+ 6" means 6 is being added to the fraction.

OpenStudy (mathstudent55):

To undo the operations, you must undo them in the opposite order they were done. Also you must use the opposite operation.

OpenStudy (anonymous):

7.2

OpenStudy (mathstudent55):

The first operation is the fraction. It is divide by 5. To undo it we do the opposite operation which is multiply by 5. The second operation is add 6. To undo it we subtract 6

OpenStudy (mathstudent55):

To undo the operations, we must undo them in THE OPPOSITE ORDER they were done. That means we start by undoing the addition of 6 by subtracting 6.

OpenStudy (anonymous):

36/5

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

i need hep

OpenStudy (mathstudent55):

Sorry. I lost connection. It always happens when I am on OS for a while.

OpenStudy (mathstudent55):

When you deal with an equation, you must do the same operation to both sides of the equation. Since we need to start by subtracting 6, we subtract 6 from both sides.